Solve the system:
x-y-z=1
-x-y+2z=-2
-x-3y+z=-5

Answer by Alan3354(69443)

Solve the system:
x-y-z=1
-x-y+2z=-2
-x-3y+z=-5
Using determinants:
X Y Z

1 -1 -1 1
-1 -1 2 -2
-1 -3 1 -5

Det = 4

X = 3
Y = 1
Z = 1
Other methods can be used, but I have an Excel sheet that does these using determinants. If you want another method, or more info on the Excel sheet, email me at gsihoutx@aol.com